# Moonshot: Optimizing Chain-Based Rotating Leader BFT via Optimistic Proposals

### v1

@@ Abstract
Existing chain-based rotating leader BFT SMR protocols for the partially synchronous network model that commit blocks with $O(1)$ minimum latency have block periods of at least $2\delta$ (where $\delta$ is the message transmission latency). While a protocol with a block period of $\delta$ exists under the synchronous model, its minimum commit latency is linear in the size of the system.
To close this gap, we present the first chain-based BFT SMR protocols with best-case delays of $\delta$ between the proposals of distinct honest leaders, and minimum commit latencies of $3\delta$. We present three protocols for the partially synchronous network model under different notions of optimistic responsiveness, two of which implement pipelining and one of which does not. All of our protocols achieve reorg resilience and two have short view lengths; properties that many existing chain-based BFT SMR protocols lack. We experimentally evaluate our protocols and show that they achieve significant increases in throughput and reductions in latency compared to the state-of-the-art, Jolteon. Our results also demonstrate that techniques commonly employed to reduce communication complexity—such as vote-pipelining and the use of designated vote-aggregators—actually reduce practical performance in many settings.
@@ I Introduction
Blockchain networks have become increasingly popular as mechanisms for facilitating decentralised, immutable and verifiable computation and storage. These networks leverage Byzantine fault-tolerant (BFT) consensus protocols to ensure that their participants (called nodes) execute the same sequence of operations (called transactions), despite some of them exhibiting arbitrary failures. Many blockchain networks also prioritize fairness; i.e., they strive to ensure that i) client transactions are processed promptly, without granting any client an unfair advantage over the others, and ii) nodes have an equal opportunity to be rewarded for the work that they do in the system. Public blockchain networks in particular also tend to be large, supporting hundreds (e.g. [24]) or thousands (e.g. [6]) of nodes in the pursuit of decentralization, and aim to cater to many concurrent clients. Accordingly, the consensus protocols driving these networks need to be efficient, maximizing transaction throughput and minimizing end-to-end commit latency (i.e., the time between a client submitting a transaction and it being executed by the blockchain).
@@
To these ends, prior works [30, 34, 18, 4, 11, 19, 25] have leveraged two key strategies: i) block chaining, and; ii) frequent leader rotation. In the block chaining (or chained) paradigm, transactions are grouped into blocks that explicitly reference one or more existing blocks (called the parents of the block), typically by including their hashes. This enables an optimization called pipelining, wherein the vote acknowledgement messages sent by the nodes in the course of agreeing upon one block can be counted towards the finalization of its parents, reducing the communication and computational complexity of the protocol by a constant factor. Our work focuses on the chain-based subcategory of chained protocols, wherein each block has exactly one parent, as opposed to DAG-based protocols in which a block may have many parents. In rotating-leader chain-based protocols, the leader responsible for proposing these blocks is changed at regular intervals, even when functioning correctly. This helps to fairly distribute the proposal workload and any related rewards. Additionally, the more frequently leaders are rotated the less amount of time a Byzantine (faulty) leader has to manipulate the ordering of pending transactions, improving censorship resistance. Accordingly, rotating-leader protocols often rotate the leader after every block proposal, an approach called Leader-Speaks-Once (LSO). This paper seeks to optimize chain-based BFT consensus performance in a modified version of the LSO setting, which we name Leader-Certifies-One (LCO). Whereas an LSO protocol allows a leader to propose only a single block, an LCO protocol allows it to propose multiple but ensures that it produces no more than one certified block during its tenure. Even as the previously cited works need not be implemented as LSO, our protocols need not be implemented as LCO, however, it is in this setting that they have the greatest advantage. Going forward, we refer to chain-based BFT consensus protocols that implement leader rotation as Chain-based Rotating Leader (CRL) protocols, and to their specific LSO or LCO variants as CLSO protocols and CLCO protocols, respectively.
@@
TABLE I: Theoretical comparison of chained rotating leader BFT SMR protocols, after GST
@@
Model Minimum Commit Latency Minimum View Change Block Period Reorg Resilience View Length Pipelined Communication Complexity Optimistic Responsiveness
@@
Simplex [13] psync. $3\delta$ $2\delta$ $3\Delta$ ${}^{+}$ $O(n^{2})$${}^{**}$
@@
*$7\delta$ assumes that the next leader aggregates votes for the current leader’s proposal; in the original HotStuff specification, leaders aggregate their own votes and forward the resulting QC to the next leader, incurring an additional $3\delta$. ${}^{+}$implex [13] requires the entire committed blockchain to be included in each proposal. ** Simplex [13] claims responsiveness only when all nodes are honest.
@@
A recent line of works [11, 13] designed CRL protocols with minimum commit latencies of $3\delta$. However, these protocols are in the non-pipelined setting, have minimum view change block periods of $2\delta$ and either have long view lengths [11] or are less practical in nature [13]. To the best of our knowledge, Apollo [5] is the only existing CRL protocol with a minimum view change block period of $\delta$. However, it incurs a minimum commit latency of $(f+1)\delta$ even during failure-free executions and assumes a synchronous network. As far as we know, no chain-based consensus protocol has simultaneously achieved a minimum view change block period of $\delta$ and a constant commit latency. To close this gap, our paper explores the design of such protocols, which we collectively refer to as Moonshot protocols.
I-A Contributions and Paper Structure
@@
We first present two state machine replication (SMR) protocols for the pipelined setting (Sections III and IV), each of which satisfies a different notion of responsiveness: i) optimistic responsiveness [34] (Definition 6) and ii) optimistic responsiveness under consecutive honest leaders [19] (Definition 7). Informally, the former requires an honest leader to make progress in $O(\delta)$ time after GST (i.e., without waiting for $\Omega(\Delta)$ time) while the latter requires an honest leader to make progress in $O(\delta)$ time only when the previous leader is also honest. Our first protocol satisfies the latter definition and is simpler to reason about, but has a longer view length. The second satisfies the former definition and has a shorter view length, but is more complex.
Both of our protocols require only two consecutive honest leaders after GST to commit a new block, and achieve reorg resilience through vote-multicasting. This strategy, together with an optimization that we call optimistic proposal, also enables them to achieve both a minimum view change block period of $\delta$ and a minimum commit latency of $3\delta$. We say that a protocol implements optimistic proposal if a leader is allowed to “optimistically” extend a block proposed by its predecessor without waiting to observe its certification. We implement this in our protocols by allowing the leader of the next view to propose a new block when it votes for a block made by the leader of the current view. Accordingly, we obtain the following result:
Theorem 1.
There exists a pipelined CRL protocol tolerating $f<n/3$ Byzantine faults with a minimum view change block period of $\delta$ and a minimum commit latency of $3\delta$.
@@
As we previously mentioned, pipelining reduces the communication and computational overhead of a CRL protocol by a constant factor. However, although this gives pipelined protocols good latency when all messages require a similar amount of time to propagate and process, pipelining actually increases commit latency when blocks take sufficiently longer to propagate or process than votes. Accordingly, we also present a non-pipelined variant of our second protocol in Section V, giving us the following result:
Theorem 2.
There exists a non-pipelined CRL protocol tolerating $f<n/3$ Byzantine faults with a minimum view change block period of $\delta$ and a minimum commit latency of $3\delta$.
This final protocol retains standard optimistic responsiveness and requires only a single honest leader to commit a new block after GST.
@@
[…]
Related work.
Finally, we present a more detailed discussion of related work in Section VII.
@@ II Preliminaries
We consider a system comprised of a set $\mathcal{V}=($$P_{1}$, …, $P_{n}$ )))) of $n$ nodes running a protocol $\mathcal{P}$ in a reliable, authenticated all-to-all network. We assume the existence of a static adversary that may corrupt up to $f<n/3$ of the nodes when $\mathcal{P}$ begins, which it may then cause to behave arbitrarily. We refer to all nodes under the control of the adversary as being Byzantine, while we refer to those that adhere to $\mathcal{P}$ as being honest. We define a quorum as a set of $\lceil\frac{n+f+1}{2}\rceil$ nodes. Henceforth, for the sake of simplicity, we assume that $n=3f+1$ and that a quorum therefore contains $12f+$ 1 nodes.
We consider the partial synchrony model of Dwork et al. [17]. Under this model, the network starts in an initial state of asynchrony during which the adversary may arbitrarily delay messages sent by honest nodes. However, after an unknown time called the Global Stabilization Time (GST), the adversary must ensure that all messages sent by honest nodes are delivered to their intended recipients within $\Delta$ time of being sent. In our initial analyses, we use $\delta$ to characterize the actual (variable) transmission latencies of messages of all types after GST and observe that $\delta\leq\Delta$. However, in our later analyses we update our model to the modified partially synchronous model of Blum et al. [7], wherein small messages (such as votes) are delivered within $\rho$ time after GST while large messages (such as block proposals) are delivered within $\Lambda$ time such that $\rho<\Lambda$.
@@ II-A Property Definitions
We note that in [34], the term “make progress” means that all honest nodes will vote for the correct (honest) leader’s proposal, not that all honest nodes observe a certificate for the included block; i.e. optimistic responsiveness does not imply reorg resilience. We also clarify that for LSO protocols, these (at most) $n-f$ responses should be messages from the previous view.
@@
We say that a CRL protocol has optimistically responsive (consecutive honest) leader change if, after GST, for any two consecutive honest leaders $L_{i}$ and $L_{i+1}$, $L_{i+1}$ sends its input within $O(\delta)$ time after receiving $L_{i}$’s input.
Importantly, this second variant of optimistic responsiveness allows the protocol to wait for $\Omega(\Delta)$ time before proposing in the new view when the leader of the previous view is Byzantine.
@@ II-B Protocol Definitions
In our protocols, a node sends a signed $\mathsf{vote}$ message to indicate its acceptance of a block. A block certificate $\mathcal{C}_{v}(B_{k})$ for view $v$ consists of a quorum of distinct signed $\mathsf{vote}$ messages for $B_{k}$ for $v$. We use $\mathcal{C}_{v}$ to denote a block certificate for view $v$ when knowledge of the related block is irrelevant to the context. We rank block certificates by their view such that $\mathcal{C}_{v}\leq\mathcal{C}_{v^{\prime}}$ if $v\leq v^{\prime}$. We provide more detailed definitions in the following sections where necessary.
@@ III Simple Moonshot
[t]
A Simple Moonshot node $P_{i}$ runs the following protocol whilst in view $v$:
$P_{i}$ additionally performs the following action in any view:
@@ A Simple Moonshot node $P_{i}$ runs the following protocol whilst in view $v$:
Prior CRL protocols require $L_{v}$ to observe $C_{v-1}$ before proposing during their happy paths (i.e. when views progress without any honest node sending a timeout message—as opposed to the fallback path). This is intended to help honest leaders create blocks that will become committed, but is unnecessarily strict for this purpose and naturally affects $\omega$ $\geq 2\delta$ and $\lambda$ $\geq 4\delta$ in the pipelined setting. Our protocols improve upon these results by requiring i) the leader of view $v$ to propose a block for $v$, say $B_{k}$, upon voting for a block in $v-1$, say $B_{k-1}$, and; ii) nodes to multicast their votes. Allowing leaders to propose optimistically in this way enables voting for $B_{k-1}$ to proceed in parallel with the proposal of $B_{k}$. Moreover, when the dissemination times of vote and proposal messages are equal, having nodes multicast their votes ensures that if all honest nodes vote for $B_{k-1}$ then they will all receive $B_{k}$ at the time that they construct $\mathcal{C}_{v-1}(B_{k-1})$, allowing them to vote for $B_{k}$ and $L_{v+1}$ to propose immediately upon entering $v$. Hence, in the happy path, $L_{v+1}$ proposes as soon as it receives $L_{v}$’s proposal, giving our protocols an $\omega$ of $\delta$. Furthermore, since our pipelined protocols require two consecutive views to produce certified blocks before a new block can be committed, requirements (i) and (ii) also give our protocols a $\lambda$ of $3\delta$.

### v2

@@ Abstract
Existing chain-based rotating-leader BFT SMR protocols for the partially synchronous network model with constant commit latencies incur block periods of at least $2\delta$ (where $\delta$ is the message transmission latency). While a protocol with a block period of $\delta$ exists under the synchronous model, its commit latency is linear in the size of the system.
To close this gap, we present the first chain-based BFT SMR protocols with $\delta$ delay between the proposals of consecutive honest leaders and commit latencies of $3\delta$. We present three protocols for the partially synchronous model under different notions of optimistic responsiveness, two of which implement pipelining. All of our protocols achieve reorg resilience and two have short view lengths; properties that many existing chain-based BFT SMR protocols lack. We present an evaluation of our protocols in a wide-area network wherein they demonstrate significant increases in throughput and reductions in latency compared to the state-of-the-art, Jolteon. Our results also demonstrate that techniques commonly employed to reduce communication complexity—such as vote-pipelining and the use of designated vote-aggregators—actually reduce practical performance in many settings.
@@ I Introduction
Blockchain networks have become increasingly popular as mechanisms for facilitating decentralised, immutable and verifiable computation and storage. These networks leverage Byzantine fault-tolerant (BFT) consensus protocols to ensure that their participants (called nodes) execute the same sequence of operations (called transactions), despite some of them exhibiting arbitrary failures. Many blockchain networks also prioritize fairness; i.e., they strive to ensure that i) client transactions are processed promptly, without granting any client an unfair advantage over the others, and ii) nodes have an equal opportunity to be rewarded for the work that they do in the system. Public blockchain networks in particular also tend to be large, supporting hundreds (e.g. [26]) or thousands (e.g. [6]) of nodes in the pursuit of decentralization, and aim to cater to many concurrent clients. Accordingly, the consensus protocols driving these networks need to be efficient, maximising transaction throughput and minimising end-to-end commit latency (i.e., the time between a client submitting a transaction and it being executed by the blockchain).
@@
To these ends, prior works [32, 37, 20, 4, 11, 21, 27] have leveraged two key strategies: i) block chaining, and; ii) frequent leader rotation. In the block chaining (or chained) paradigm, transactions are grouped into blocks that explicitly reference one or more existing blocks (called the parents of the block), typically by including their hashes. This enables an optimization called pipelining, wherein the vote acknowledgement messages sent by the nodes in the course of agreeing upon a block can be counted towards the finalization of its parents, facilitating the removal of additional voting phases and thus reducing the communication and computational complexity of the protocol by a constant factor. Our work focuses on the chain-based subcategory of chained protocols, wherein each block has exactly one parent, as opposed to DAG-based protocols in which a block may have many parents. In rotating-leader chain-based protocols, the leader responsible for proposing these blocks is changed at regular intervals, even when functioning correctly. This helps to fairly distribute the proposal workload and any related rewards. Additionally, the more frequently leaders are rotated the less amount of time a Byzantine (faulty) leader has to manipulate the ordering of pending transactions, improving censorship resistance. Accordingly, rotating-leader protocols often rotate the leader after every block proposal, an approach called leader-speaks-once (LSO). This paper seeks to optimize chain-based BFT consensus performance in a modified version of the LSO setting, which we name leader-certifies-one (LCO). Whereas an LSO protocol allows a leader to propose only a single block, an LCO protocol allows it to propose multiple but ensures that it produces no more than one certified block. Even as the previously cited works need not be implemented as LSO, our protocols need not be implemented as LCO, however, it is in this setting that they have the greatest advantage. We henceforth refer to chain-based BFT consensus protocols that implement leader rotation as CRL protocols.
@@
TABLE I: Theoretical comparison of chain-based rotating leader BFT SMR protocols
@@
Model Minimum Commit Latency Minimum View Change Block Period Reorg Resilience View Length Pipelined Communication Complexity(1) Optimistic Responsiveness
@@
Simplex [13] psync. $3\delta$ $2\delta$ $3\Delta$ Unbounded(3) $O(n^{2})$ (4)
@@
(1) Assuming the use of threshold signatures. (2) HotStuff has a minimum commit latency of $7\delta$ if the next leader aggregates the votes for the current leader’s proposal. In the original HotStuff specification, leaders aggregate the votes for their own proposals and then forward the resultant QC to the next leader, incurring an additional $3\delta$. (3) Simplex [13] requires each proposal to include its notarized parent blockchain, making the size of each proposal proportional to the size of the blockchain itself. (4) Simplex [13] claims responsiveness only when all nodes are honest.
@@
A recent line of work [11, 13] designed CRL protocols with minimum commit latencies of $3\delta$. However, these protocols are in the non-pipelined setting, have minimum view change block periods of $2\delta$ and either have long view lengths [11] or are less practical in nature [13]. To the best of our knowledge, Apollo [5] is the only existing CRL protocol with a minimum view change block period of $\delta$. However, it incurs a minimum commit latency of $(f+1)\delta$ even during failure-free executions and assumes a synchronous network. As far as we know, no chain-based consensus protocol has simultaneously achieved a minimum view change block period of $\delta$ and a constant commit latency. To close this gap, our paper explores the design of such protocols, which we collectively refer to as Moonshot protocols.
I-A Contributions
@@
We first present two state machine replication (SMR) protocols for the pipelined setting, each of which satisfies a different notion of responsiveness: i) optimistic responsiveness [37] (Definition 6) and ii) optimistic responsiveness under consecutive honest leaders [21] (Definition 7). Informally, the former requires an honest leader to make progress in $O(\delta)$ time after GST (i.e., without waiting for $\Omega(\Delta)$ time) while the latter requires an honest leader to make progress in $O(\delta)$ time only when the previous leader is also honest. Our first protocol satisfies the former definition and is simpler to reason about, but has a longer view length. The second satisfies the latter definition and has a shorter view length, but is more complex.
Both of our protocols require only two consecutive honest leaders after GST to commit a new block, and achieve reorg resilience through vote-multicasting. This strategy, together with an optimization that we call optimistic proposal, also enables them to achieve both a minimum view change block period of $\delta$ and a minimum commit latency of $3\delta$. We say that a protocol implements optimistic proposal if a leader is allowed to “optimistically” extend a block proposed by its predecessor without waiting to observe its certification. We implement this in our protocols by allowing the leader of the next view to propose a new block when it votes for a block made by the leader of the current view.
@@
As mentioned, pipelining can reduce the communication and computational overhead of a protocol. However, while this gives pipelined protocols good latency when all messages require a similar amount of time to propagate and process, pipelining actually increases commit latency when blocks take sufficiently longer to propagate or process than votes. Accordingly, we also present a non-pipelined variant of our second protocol in Section V. This final protocol retains standard optimistic responsiveness and requires only a single honest leader to commit a new block after GST.
@@
Subsequently, we present an evaluation of LCO implementations of our protocols against an LSO implementation of Jolteon. Our protocols outperformed Jolteon in failure-free wide-area networks (WANs) of up to $200$ nodes, committing approximately $1.5$x as many blocks at around half the latency, on average. Our protocols also outperformed under failures, with our non-pipelined protocol committing $8$x as many blocks with a reduction in latency by more than two orders of magnitude under Jolteon’s worst-case leader schedule.
Organization.
The rest of the paper is organized as follows: In Section II, we present the system model and preliminaries for our work. Section III presents a pipelined CRL protocol with a minimum commit latency of $3\delta$, minimum view change block period of $\delta$, reorg resilience and optimistic responsiveness under consecutive honest leaders. We then modify this protocol in Section IV to obtain a protocol with standard optimistic responsiveness and improved view length. In Section V, we give a non-pipelined version of our second protocol, which offers improved commit latency when blocks take sufficiently longer to propagate or process than votes. We present an evaluation of our protocols in Section VI, and conclude with a more detailed discussion of related works in Section VII.
@@ II Preliminaries
We consider a system comprised of a set $\mathcal{V}=(P_{1},\ldots,P_{n})$ of $n$ nodes running a protocol $\mathcal{P}$ in a reliable, authenticated all-to-all network. We assume the existence of a static, computationally bounded adversary that cannot break cryptographic primitives but may corrupt up to $f<n/3$ of the nodes when $\mathcal{P}$ begins, which it may then cause to behave arbitrarily. We refer to all nodes under the control of the adversary as being Byzantine, while we refer to those that adhere to $\mathcal{P}$ as being honest. We define a quorum as a set of $\lfloor\frac{n}{2}\rfloor+f+1$ nodes. Henceforth, for the sake of simplicity, we assume that $n=3f+1$ and that a quorum therefore contains $12f+$ 1 nodes.
We assume that each node has access to a local clock and that these clocks collectively have no drift and arbitrary skew. We also assume the partially synchronous communication model of Dwork et al. [18]. Under this model, the network starts in an initial state of asynchrony during which the adversary may arbitrarily delay messages sent by honest nodes. However, after an unknown time called the Global Stabilization Time (GST), the adversary must ensure that all messages exchanged between honest nodes are delivered within $\Delta$ time of being sent (from the perspective of the sender). In our initial analyses, we denote the range of the actual transmission latencies of messages of all types with $\delta$, and observe that $\delta=[0,\Delta]$ after GST. Moreover, when we measure latency in terms of $\delta$, e.g. $x=y\delta$, we are denoting that $x$ requires the propagation of $y$ sequential messages (i.e. $x$ requires $y$ network hops). In our later analyses we base our communication model on the modified partially synchronous model of Blum et al. [7]. Under this model, we denote the range of the actual delivery times of small messages (such as votes) with $\rho$ and that of large messages (such as block proposals) with $\beta$, such that $\rho=[0,\min(\beta))$ and $\beta=(\max(\rho),\Delta]$, after GST. We follow a similar convention as with $\delta$ when measuring latency in terms of $\beta$ and $\rho$, with $x=y\beta+z\rho$ denoting that $x$ requires the sequential propagation of $y$ large and $z$ small messages.
@@ II-A Property Definitions
We note that in [37], the term “make progress” means that all honest nodes will vote for the correct (honest) leader’s proposal, not that all honest nodes observe a certificate for the included block; i.e. optimistic responsiveness does not imply reorg resilience. We also clarify that for LSO/LCO protocols, these (at most) $n-f$ responses should be messages from the previous view.
@@
We say that a protocol is optimistically responsive (consecutive honest) if after GST, for any two consecutive honest leaders $L_{v}$ and $L_{v+1}$, $L_{v+1}$ sends its proposal within $O(\delta)$ time of receiving $L_{v}$’s proposal.
Importantly, this variant of optimistic responsiveness allows the protocol to wait for $\Omega(\Delta)$ time before proposing in the new view when the leader of the previous view is Byzantine.
@@ II-B Protocol Definitions
In our protocols, a node sends a signed $\mathsf{vote}$ message to indicate its acceptance of a block. A block certificate $\mathcal{C}_{v}(B_{k})$ for view $v$ consists of a quorum of distinct signed $\mathsf{vote}$ messages for $B_{k}$ for $v$. We use $\mathcal{C}_{v}$ to denote a block certificate for view $v$ when knowledge of the related block is irrelevant to the context. We rank block certificates by their views such that $\mathcal{C}_{v}\leq\mathcal{C}_{v^{\prime}}$ if $v\leq v^{\prime}$. We provide more detailed definitions in the following sections where necessary.
@@ III Simple Moonshot
A Simple Moonshot node $P_{i}$ runs the following protocol whilst in view $v$: 1. Propose. If $P_{i}$ is $L_{v}$ and enters $v$ at time $t_{i}$, propose: (i) upon receiving $\mathcal{C}_{v-1}(B_{k-1})$ before $t_{i}+2\Delta$, or; (ii) at $t_{i}+2\Delta$. Do so by multicasting $\langle\mathsf{propose},B_{k},\mathcal{C}_{v^{\prime}}(B_{k-1}),v\rangle$, where $\mathcal{C}_{v^{\prime}}(B_{k-1})$ is the highest ranked block certificate known to $L_{v}$ and $B_{k}$ extends $B_{k-1}$. 2. Vote. $P_{i}$ votes once using one of the following rules: (a) Upon receiving $\langle{\sf opt\text{-}propose},B_{k},v\rangle$ such that $B_{k}$ extends $B_{k-1}$, if ${\sf lock}_{i}=\mathcal{C}_{v-1}(B_{k-1})$ then multicast $\langle\mathsf{vote},H(B_{k}),v\rangle_{i}$. (b) Upon receiving $\langle\mathsf{propose},B_{k},\mathcal{C}_{v^{\prime}}(B_{h}),v\rangle$, if $\mathcal{C}_{v^{\prime}}(B_{h})\geq{\sf lock}_{i}$ and $B_{k}$ extends $B_{h}$ then multicast $\langle\mathsf{vote},H(B_{k}),v\rangle_{i}$. 3. Optimistic Propose. Upon voting for $B_{k}$ in $v$, if $P_{i}$ is $L_{v+1}$, multicast $\langle{\sf opt\text{-}propose},B_{k+1},v+1\rangle$ such that $B_{k+1}$ extends $B_{k}$. 4. Timeout. Upon receiving $f+1$ distinct $\langle\mathsf{timeout},v\rangle_{*}$ or when ${\sf view\text{-}timer}_{i}$ expires, stop voting in $v$ and multicast $\langle\mathsf{timeout},v\rangle_{i}$. 5. Advance View. Upon receiving $\mathcal{C}_{v^{\prime}-1}(B_{h})$ or $\mathcal{TC}_{v^{\prime}-1}$, where $v^{\prime}>v$, and before executing any other rule, do the following: i) multicast the certificate; ii) update ${\sf lock}_{i}$ to the highest ranked block certificate received so far; iii) unicast a status message $\langle\mathsf{status},v^{\prime},{\sf lock}_{i}\rangle$ to $L_{v^{\prime}}$ if ${\sf lock}_{i}$ has a view less than $v^{\prime}-1$, iv) enter $v^{\prime}$, and; v) reset ${\sf view\text{-}timer}_{i}$ to $5\Delta$ and start counting down. $P_{i}$ additionally performs the following action in any view: 1. Direct Commit. Upon receiving $\mathcal{C}_{v-1}(B_{k-1})$ and $\mathcal{C}_{v}(B_{k})$ such that $B_{k}$ extends $B_{k-1}$, commit $B_{k-1}$. 2. Indirect Commit. Upon directly committing $B_{k-1}$, commit all of its uncommitted ancestors.
@@
Prior CRL protocols require $L_{v}$ to observe $C_{v-1}$ before proposing during their happy paths (i.e. when views progress without any honest node sending a timeout message—as opposed to the fallback path). This is intended to help honest leaders create blocks that will become committed, but is unnecessarily strict for this purpose and naturally affects $\omega$ $\geq 2\delta$ and $\lambda$ $\geq 4\delta$ in the pipelined setting. Our protocols improve upon these results by requiring i) the leader of view $v$ to propose a block for $v$, say $B_{k}$, upon voting for a block in $v-1$, say $B_{k-1}$, and; ii) nodes to multicast their votes. Allowing leaders to propose optimistically in this way enables voting for $B_{k-1}$ to proceed in parallel with the proposal of $B_{k}$. Moreover, when the dissemination times of vote and proposal messages are equal (see Figure 2), having nodes multicast their votes ensures that if all honest nodes vote for $B_{k-1}$ then they will all receive $B_{k}$ at the time that they construct $\mathcal{C}_{v-1}(B_{k-1})$, allowing them to vote for $B_{k}$ and $L_{v+1}$ to propose immediately upon entering $v$. Hence, in the happy path, $L_{v+1}$ proposes as soon as it receives $L_{v}$’s proposal, giving our protocols an $\omega$ of $\delta$. Furthermore, since our pipelined protocols require two consecutive views to produce certified blocks before a new block can be committed, requirements (i) and (ii) also give our protocols a $\lambda$ of $3\delta$.
Figure 2: Optimistic proposal (pictured in blue) and vote multicasting (pictured in orange) enable Simple Moonshot and Pipelined Moonshot to propose new blocks at the same rate that they become certified when proposals and votes take equal time to propagate and process.